# New polarimetric study of the galactic X-ray burster GX 13$+$1

### v1

@@ 3.2 Polarimetry
Spectroscopy revealed that neither the light curve nor the hardness have strong dips or features that would suggest separating the observation into several parts to be studied independently. To study the time variability of the polarimetric properties, we followed the same approach as in B24 and split this new IXPE observation into five equal time bins spanning 9.5 h. The results are reported in Fig. 4.
Figure 4: Time dependence of the normalized Stokes parameters $q$ and $u$ (top), and of the PD and PA (bottom) obtained by the pcube algorithm using the data separated into 9.5 h time bins. The confidence regions are reported at $68\%$.
Figure 5: Energy resolved normalized Stokes parameters $q$ and $u$ (top), and PD and PA (bottom) obtained by the pcube algorithm using the reported energy bins for the time-averaged data. The confidence regions are reported at $68\%$ CL. The vectors correspond to the best-fit polarization parameters of the diskbb and bbodyrad components, respectively, as obtained in the spectro-polarimetric analysis and reported in Table 4.
@@ 3.3 Spectro-polarimetry
Table 4: IXPE spectro-polarimetric fit of the Stokes parameters $I$, $Q$, and $U$ using the spectral model from Table 3.
Model PD/$A_{1}$ (%) $A_{\rm slope}$ (keV-1) PA/$\psi_{1}$ (deg) $\psi_{\rm slope}$ (deg,keV-1) $\chi^{2}$/d.o.f.
@@
Figure 6: Spectral energy distribution of GX 13$+$1 in $EF_{E}$ representation as observed by IXPE. The left, middle, and right panels are for the Stokes parameters $I$, $Q$, and $U$, respectively. The fit is performed in the 2–8 keV energy band using the three IXPE detectors and applying the model tbabs*(diskbb*polconst+bbodyrad*polconst). The total model is shown with the solid black line, while the diskbb and bbodyrad with the dashed and dotted lines, respectively. The lower sub-panels show the residuals between the data and the best fit. In the joint spectral fit, we apply a rebinning to have at least 50 counts per bin as in Swift/XRT, while for the spectro-polarimetric analysis, a constant grouping is applied to have energy bins of 200 eV.
@@ 4 Discussion
Although interesting as a single observation, the results reported in this article provide the most information when compared to the results of the October observations reported in B24. Overall, in October IXPE measured a PD of $\sim$1.4and a PA of $\sim-2\degr$, significantly different from the current results. However, the reason for the depolarization in the first observation was fast variability in the PA. In B24, the observation was separated into the pre-dip, dip, and post-dip states. The results for the dip were poorly constrained, but the other two are presented in Fig. 7 and compared to the results of the current observation. We note that the current observation is similar in PA to the post-dip state of B24. The lower PD in the entire current observation most probably comes from averaging the entire observation in time and energy, as for the previous one. In Fig. 4, the Bin 3, for instance, shows a polarization within errors similar to the one measured after the dip in B24.
@@
Both HID and the shape of the spectrum confirm that GX 13$+$1 was in a soft state throughout the observation. We do not see any significant change of state in the IXPE and Swift/XRT energy bands, although, in the last part of the observation, the flux rises, which is a sign of the source going into an even softer state. If we look at the HID of the source in the October observation (B24, Figure 2), we see that the source is almost in the same state, except for the dipping period. In this new observation, since simultaneous Swift/XRT observations offer a better spectral capability than IXPE, the spectral modeling and decomposition into two main components are better determined, allowing for a better spectro-polarimetric analysis.
The light curve does not show any peculiar variability: unlike the one from the October observation, the current one has no strong dips or rapid changes, and neither does the hardness ratio. Despite this constant hardness, we attempted to study the possible variation in 9.5 h equal time bins similarly to B24. This analysis reports a hint for a variation in PD in Fig. 4, in particular is interesting that in Bin 5, corresponding to a softening of GX 13$+$1 there is a lower polarization and an indication for a possible rotation with respect to the other 4 time bins. The clear rotation of PA observed during the first observation is not present this time, but the PA appears to have remained aligned with the post-dip polarization measured in B24. If we compare the positions of the points on the $(q,u)$ plane with those obtained from the October 2023 observation (Figure 7 in B24), we see a different pattern, but the points are still in the I or IV quadrant. The biggest difference is in order: in October, we saw the points following the straight line path, but here the variability shows a different pattern.
@@
From the aligned systems, we expect the polarization vectors of the emission coming from the disk and the Comptonized component to be either nearly parallel or perpendicular to each other. Hence, we expect some misalignment in the system that could support this difference in the PA. Another thing to notice is a very high PD of the disk component: as was mentioned in Loktev et al. (2022), we do not expect a polarization higher than 4% in the IXPE energy range even for the highest inclination. This value is compatible with the spectro-polarimetric results at 90% CL. However, knowing that the source is likely to have an inclination lower than 80°(Díaz Trigo et al., 2012; Tomaru et al., 2020), we need to add the scattering of the disk emission in the wind to reach the PD of 6.5%. We assume that the Comptonized component comes from the boundary layer, as it is expected to align in PA with the disk emission, and the difference in PAs is less than 45°.
@@
In search of a source of misalignment, we looked at the previous observations of IXPE. It is possible (although unexpected) that, similar to the Cir X-1 (Rankin et al., 2024), the rotation axis of the neutron star is slightly misaligned from the rotation axis of the binary system. The boundary layer would then be dragged out of the disk plane by the motion of the star. Another option would be the peculiar geometry of the wind in the system.
@@
Finally, we want to address the significant difference between the two observations of GX 13$+$1. The object was in the same spectral state but half an orbit away from the previous observation, when the source clearly showed a rotation of the PA with time. However, in February 2024, there was no clear monotonic change in the PA, while we still see variations in the PA. One possibility is that the mechanism responsible for the dip in the light curve during the first observation was also responsible for the rotation of the PA. However, this mechanism is unknown. What we can address is that from this new observation, it appears that the PA remained aligned with the one observed in the end of the previous observation. Then, at the end of the present IXPE pointing, when the flux grows and the source goes into a softer state, the PD decreases and rotates to a PA similar to the one we observed in October before the dip. This could suggest, similarly to what we observed in Cir X-1, a variation of the PA with the state of the source. This could mean a variation of the geometry of the hot region in different states; this can also be confirmed by the indication we obtained in the October 2023 observation when the energy dependence of the polarization changed before and after the dip going from an energy-dependent behavior, as observed e.g. for 4U 1820--30 (Di Marco et al., 2023a), into a constant polarization with energy, as observed, e.g., for Sco X-1 in La Monaca et al. (2024). Present data do not allow for a clear statement on this, but we hope in the future we could obtain further observations to improve the significance for a detailed analysis of the polarization along the states of atoll and Z sources.
A further point is the contribution of the wind, as we are unaware of the wind behavior during both observations, we can also assume that it could be responsible for the differences in polarimetry, yet we have no way to support this claim.
@@ 5 Summary
We analyzed the observation of GX 13$+$1 carried out simultaneously by IXPE and Swift/XRT. In this article, we report the highly significant detection of polarization from this source in the soft state. Spectroscopic analysis showed that the spectrum is close to that in Díaz Trigo et al. (2012), and the hardness of the emission confirmed the lower left banana state, as well as the further softening of the spectrum at the end of this new observation. We used a diskbb model for the softer disk emission, and the bbodyrad model to approximate the harder component associated with the spreading or boundary layer. The polarimetric analysis provided the overall PD of the source during the observation at 2.5% at ¿$5\sigma$ CL, and a PA at 24°. We also saw a marginally significant decrease in PD and PA with energy. The same trend was confirmed in the spectro-polarimetric analysis, which showed that the softer component is stronger polarized than the harder component with the PA differing by $\sim$40°. The results of the spectro-polarimetric analysis using xspec performed under the assumption of one overall polarization model agreed with the pcube polarimetric results.
We studied the differences and similarities between the first (B24) and the second observations of GX 13$+$1 by IXPE. Our main conclusion is that we saw a significantly higher overall polarization in the second observation, even though the source was in the same state during both observations. The depolarization comes from the variability of polarimetric properties, which was larger in the first observation. We discussed possible reasons for the different polarimetric behavior in the two observations.
@@ Acknowledgements.
This research used data products provided by the IXPE Team (MSFC, SSDC, INAF, and INFN) and distributed with additional software tools by the High-Energy Astrophysics Science Archive Research Center (HEASARC), at NASA Goddard Space Flight Center (GSFC). The Imaging X-ray Polarimetry Explorer (IXPE) is a joint US and Italian mission. The Italian contribution is supported by the Italian Space Agency (Agenzia Spaziale Italiana, ASI) through contract ASI-OHBI-2022-13-I.0, agreements ASI-INAF-2022-19-HH.0 and ASI-INFN-2017.13-H0, and its Space Science Data Center (SSDC) with agreements ASI-INAF-2022-14-HH.0 and ASI-INFN 2021-43-HH.0, and by the Istituto Nazionale di Astrofisica (INAF) and the Istituto Nazionale di Fisica Nucleare (INFN) in Italy. We thank the Swift Project Scientists for approving our DDT request to observe GX 13$+$1. This research has been supported by the Finnish Cultural Foundation grant 00240328 (AB) and the Academy of Finland grant 333112 (AB, JP, SVF, VL). ADM, FLM are partially supported by MAECI with grant CN24GR08 “GRBAXP: Guangxi-Rome Bilateral Agreement for X-ray Polarimetry in Astrophysics”.

### v2

@@ 3.2 Polarimetry
The spectroscopy reveals that neither the light curve nor the hardness has strong dips or features that would suggest the observation should be separated into several parts to be studied independently. To study the time variability of the polarimetric properties, we followed the same approach as in B24 and split this new IXPE observation into five equal time bins spanning 9.5 h. The results are reported in Fig. 4.
Figure 4: Time dependence of the normalized Stokes parameters $q$ and $u$ (top), and of the PD and PA (bottom) obtained by the pcube algorithm using the data separated into 9.5 h time bins. Confidence regions are reported at $68\%$.
Figure 5: Energy-resolved normalized Stokes parameters $q$ and $u$ (top), and PD and PA (bottom) obtained by the pcube algorithm using the reported energy bins for time-averaged data. The confidence regions are reported at $68\%$ CL. The vectors correspond to the best-fit polarization parameters of the diskbb and bbodyrad components, respectively, as obtained using spectropolarimetric analysis and reported in Table 4.
@@ 3.3 Spectropolarimetry
Table 4: IXPE spectropolarimetric fit of the Stokes parameters $I$, $Q$, and $U$ using the spectral model from Table 3.
Model PD/$A_{1}$ (%) $A_{\rm slope}$ (keV-1) PA/$\psi_{1}$ (deg) $\psi_{\rm slope}$ (deg keV-1) $\chi^{2}$/d.o.f.
@@
Figure 6: Spectral energy distribution of GX 13$+$1 in $EF_{E}$ representation as observed by IXPE. The left, middle, and right panels are for the Stokes parameters $I$, $Q$, and $U$, respectively. The fit is performed in the 2–8 keV energy band using the three IXPE detectors and applying the tbabs*(diskbb*polconst+bbodyrad*polconst) model. The total model is shown with the solid black line, while diskbb and bbodyrad are indicated with the dashed and dotted lines, respectively. The lower subpanels show the residuals between the data and the best fit. In the joint spectral fit, we applied a rebinning to obtain at least 50 counts per bin as in Swift/XRT, while for the spectropolarimetric analysis, a constant grouping was applied to have energy bins of 200 eV.
@@ 4 Discussion
Although intriguing on its own, the observation reported in this article yields the most insight when compared to the results of the October 2023 observations reported in B24, when we measured an overall PD of $\sim$1.4and a PA of $\sim-2\degr$, significantly different from the current results. However, the reason for the depolarization in the first observation was fast variability in the PA. In B24, the observation was separated into the pre-dip, dip, and post-dip states. The results for the dip were poorly constrained, but the other two are presented in Fig. 7 and compared to the results of the current observation. We note that the current observation is similar in PA to the post-dip state of B24, but the PD is lower probably due to the averaging over time and energy. Bin 3 in Fig. 4, for instance, shows a polarization within errors similar to the one measured after the dip in B24.
@@
Both the HID and the shape of the spectrum confirm that GX 13$+$1 was in a soft state throughout the observation. We do not see any significant change of state in the IXPE and Swift/XRT energy bands, although, in the last part of the observation, the flux rises, which is a sign of the source going into an even softer state. If we look at the HID of the source in the October observation (Fig. 2 in B24), we see that the source is almost in the same state, except for the dipping period. During the new observation period, the spectral modeling and decomposition into two main components are better determined, since simultaneous Swift/XRT observations offer a better spectral capability than IXPE, thereby allowing for a better spectropolarimetric analysis.
The light curve does not show any peculiar variability: unlike the one from the October observation, the current one has no strong dips or rapid changes, and neither does the hardness ratio. Despite this constant hardness, we attempted to study the possible variation in 9.5 h equal time bins similar to B24. This analysis suggests possible variation in PD in Fig. 4; in particular, it is interesting to note that in Bin 5, corresponding to a softening of GX 13$+$1, there is lower polarization and an indication of a possible rotation with respect to the other four time bins. The clear rotation of the PA observed during the first observation is not present this time, but the PA appears to have remained aligned with the post-dip polarization measured in B24. If we compare the positions of the points on the $(q,u)$ plane with those obtained from the October 2023 observation (Fig. 7 in B24), we see a different pattern; however, the points are still in the I or IV quadrant. The most significant difference lies in the pattern: in October, we observed the points following a straight line path, whereas here, the variability exhibits a different pattern.
@@
From the aligned systems, we expect the polarization vectors of the emission coming from the disk and the Comptonized component to be either nearly parallel or perpendicular to each other. Hence, we expect some misalignment in the system that would account for this difference in the PA. Additionally, it is important to note the exceptionally high PD of the disk component. As stated by Loktev et al. (2022), polarization exceeding 4% in the IXPE energy range is not anticipated, even at the highest inclination. This value is compatible with the spectropolarimetric results at 90% CL. However, knowing that the source is likely to have an inclination lower than 80°(Díaz Trigo et al., 2012; Tomaru et al., 2020), we need to add the scattering of the disk emission in the wind to reach a PD of 6.5%. We assumed that the Comptonized component comes from the boundary layer, as it is expected to align in PA with the disk emission. The observed difference in PAs is less than 45°.
@@
In search of a source of misalignment, we examined previous observations by IXPE. It is possible, though unexpected, that the rotation axis of the neutron star is slightly misaligned from the rotation axis of the binary system, similar to the Cir X-1 (Rankin et al., 2024). In such a scenario, the boundary layer could be dragged out of the disk plane by the motion of the star. Another potential explanation is the peculiar geometry of the wind within the system.
@@
Finally, we wish to address the significant difference between the two observations of GX 13$+$1. Although the object was in the same spectral state, it was observed half an orbit away from the previous observation, during which the source clearly exhibited a rotation of the PA over time. However, in February 2024, there was no evident monotonic change in the PA, despite ongoing variations. One possibility is that the mechanism responsible for the dip in the light curve during the first observation also caused the rotation of the PA. However, this mechanism remains unidentified. Notably, the recent observation suggests that the PA remained aligned with that observed at the end of the previous observation. As the flux increased and the source transitioned to a softer state at the end of the present IXPE pointing, the PD decreased and rotated to a PA similar to that observed in October before the dip. This pattern might indicate, as observed in Cir X-1, a variation in the PA corresponding to the state of the source. This implies a potential variation in the geometry of the hot region across different states. This hypothesis is supported by the October 2023 observation, where the energy dependence of the polarization shifted from an energy-dependent behavior, as seen in 4U 1820--30 (Di Marco et al., 2023), to a constant polarization with energy, as observed, for example, in Sco X-1 in La Monaca et al. (2024). While current data do not allow for a definitive conclusion, we hope that future observations will provide sufficient significance for a detailed analysis of the polarization across different states of atoll and Z sources.
Additionally, the role of the wind remains an open question. Given our lack of knowledge regarding wind behavior during both observations, it is plausible that the wind could contribute to the observed differences in polarimetric properties, although we currently have no means to substantiate this hypothesis.
@@ 5 Summary
We analyzed the observation of GX 13$+$1 carried out simultaneously by IXPE and Swift/XRT. In this article, we report the highly significant detection of polarization from this source in the soft state. Spectroscopic analysis shows that the spectrum is close to that in Díaz Trigo et al. (2012), and the hardness of the emission confirms the lower left banana state, as well as the further softening of the spectrum at the end of this new observation. We used a diskbb model for the softer disk emission, and the bbodyrad model to approximate the harder component associated with the spreading or boundary layer. The polarimetric analysis provided the overall PD of 2.5% at ¿$5\sigma$ CL and a PA of 24°. We also see a marginally significant decrease in PD and PA with energy. The same trend was confirmed by the spectropolarimetric analysis, which showed that the softer component was more strongly polarized than the harder component, with the PA differing by $\sim$40°. The results of the spectropolarimetric analysis using xspec performed under the assumption of one overall polarization model were in agreement with the pcube polarimetric results.
We studied the differences and similarities between the first (B24) and the second observations of GX 13$+$1 by IXPE. Our main conclusion is that the overall polarization is significantly higher in the second observation despite the source being in the same state during the two observations. Depolarization was caused by the variability in polarimetric properties, which was larger during the first observation. We have discussed possible reasons for the different polarimetric behavior in the two observations.
@@ Acknowledgements.
This research used data products provided by the IXPE Team (MSFC, SSDC, INAF, and INFN) and distributed with additional software tools by the High-Energy Astrophysics Science Archive Research Center (HEASARC), at NASA Goddard Space Flight Center (GSFC). The Imaging X-ray Polarimetry Explorer (IXPE) is a joint US and Italian mission. The Italian contribution is supported by the Italian Space Agency (Agenzia Spaziale Italiana, ASI) through contract ASI-OHBI-2022-13-I.0, agreements ASI-INAF-2022-19-HH.0 and ASI-INFN-2017.13-H0, and its Space Science Data Center (SSDC) with agreements ASI-INAF-2022-14-HH.0 and ASI-INFN 2021-43-HH.0, and by the Istituto Nazionale di Astrofisica (INAF) and the Istituto Nazionale di Fisica Nucleare (INFN) in Italy. We thank the Swift Project Scientists for approving our DDT request to observe GX 13$+$1. This research has been supported by the Finnish Cultural Foundation grant 00240328 (AB) and the Academy of Finland grant 333112 (AB, JP, SVF, VL). ADM and FLM are partially supported by MAECI with grant CN24GR08 “GRBAXP: Guangxi-Rome Bilateral Agreement for X-ray Polarimetry in Astrophysics”.